# Unveiling the Potential of LLM-Based ASR on Chinese Open-Source Datasets

### v2

@@ 1 Introduction
Research efforts to integrate LLMs with ASR systems generally fall into two categories. The first strategy involves connecting LLMs with pre-trained ASR models, wherein the ASR-generated text is directly fed to the LLM to serve as a prompt for downstream tasks [13, 14, 15, 16] or to facilitate error correction [17]. However, this coarse-grained integration may result in a substantial loss of acoustic information. It may suffer compounding errors from the initial ASR stage, potentially leading to an exacerbation of inaccuracies. In contrast, the second approach adopts audio-text cross-modal LLMs, which embrace the auditory modality by employing an encoder network to process the speech and generate embeddings that are subsequently provided to a decoder-only LLM [5, 18, 19, 20, 21, 22]. This framework strives for a tighter coupling between acoustic cues and linguistic context, aiming to yield a better interpretation of speech. Through a series of studies, the paradigm of augmenting a speech foundation model with an LLM through projector modules has emerged as the prevailing framework in the current LLM-based speech recognition research. Specifically, SALMONN [20] applies Whisper [23] to extract semantic content and BEATs [24] for audio event information, culminating in a robust perception of human speech, music, and audio events. Qwen-Audio [21] implements Whisper as the exclusive encoder, utilizing structured task directives to enhance the model’s performance across various audio tasks. SLAM-ASR [22] leverages a linear layer as the projector module and achieves SOTA performance on the English 960-hour LibriSpeech [25] task.
@@
We will provide reproducible recipes encompassing the entire pipeline, including data preparation, training, inference, and scoring. Furthermore, we will release pre-trained models, enabling researchers to delve into the specifics of the training process and gain valuable insights for their own investigations.
@@ 3.1 Datasets
Testing Set: To reduce bias and identify unique challenges, we test on nine open-source and two internal test sets, as shown in Table 1. Specifically, the open test sets include AISHELL-1 [29], AISHELL-2 [30], Test_Net [28], Test_Meeting [28], SPEECHIO_0, SPEECHIO_1, SPEECHIO_2, SPEECHIO_3, SPEECHIO_4222SPEECHIO_* means SPEECHIO_ASR_ZH0000* [32]. The internal noisy and accent test sets have different acoustic conditions (noisy and accented speech) from the open test sets.
@@ 4.4 Optimal configuration
We also compare our results to vanilla fine-tuned Whisper models. We use the same training set to fine-tune Whisper-large-v2, while the WeNet team[38] fine-tunes Whisper-large-v3 using WenetSpeech (corrected transcript). The experimental results in Table 6 show that our proposed model performs significantly better than the Whisper series models. This indicates that the performance could not be achieved by the speech foundation model encoder alone, and the LLM helped to improve the performance.
Table 6: Comparison with popular models in terms of CER%($\downarrow$). Whisper-L-v2 and Whisper-L-v3 have both undergone fine-tuning.
Dataset Para-former Qwen-Audio Whisper-L-v2 Whisper-L-v3 Ours
AISHELL-1 1.95 1.30 3.18 5.02 0.95
AISHELL-2 3.01 3.18 4.49 6.60 3.50
Test_Net 6.74 9.50 9.08 6.54 6.06
Test_Meeting 6.97 10.87 9.73 8.84 6.26
SPEECHIO_0 2.55 5.08 3.44 4.05 3.05
SPEECHIO_1 0.49 1.17 1.98 3.17 1.58
SPEECHIO_2 3.23 5.68 5.27 6.25 3.73
SPEECHIO_3 1.13 2.81 4.86 7.09 2.42
SPEECHIO_4 1.33 3.99 3.63 4.29 3.29

### v3

@@ 1 Introduction
Research efforts to integrate LLMs with ASR systems generally fall into two categories. The first strategy involves connecting LLMs with pre-trained ASR models, wherein the ASR-generated text is directly fed to the LLM to serve as a prompt for downstream tasks [13, 14, 15] or to facilitate error correction [16]. However, this coarse-grained integration may result in a substantial loss of acoustic information. It may suffer compounding errors from the initial ASR stage, potentially leading to an exacerbation of inaccuracies. In contrast, the second approach adopts audio-text cross-modal LLMs, which embrace the auditory modality by employing an encoder network to process the speech and generate embeddings that are subsequently provided to a decoder-only LLM [5, 17, 18, 19, 20, 21]. This framework strives for a tighter coupling between acoustic cues and linguistic context, aiming to yield a better interpretation of speech. Through a series of studies, the paradigm of augmenting a speech foundation model with an LLM through projector modules has emerged as the prevailing framework in the current LLM-based speech recognition research. Specifically, SALMONN [19] applies Whisper [22] to extract semantic content and BEATs [23] for audio event information, culminating in a robust perception of human speech, music, and audio events. Qwen-Audio [20] implements Whisper as the exclusive encoder, utilizing structured task directives to enhance the model’s performance across various audio tasks. SLAM-ASR [21] leverages a linear layer as the projector module and achieves SOTA performance on the English 960-hour LibriSpeech [24] task.
@@
We will provide reproducible recipes encompassing the entire pipeline, including data preparation, training, inference, and scoring. Furthermore, we will release pre-trained models, enabling researchers to delve into the specifics of the training process and gain valuable insights for their own investigations [^1] https://github.com/gengxuelong/wenet_LLM_from_ASLP.
@@ 3.1 Datasets
Testing Set: To reduce bias and identify unique challenges, we test on nine open-source and two internal test sets, as shown in Table 1. Specifically, the open test sets include AISHELL-1 [28], AISHELL-2 [29], Test_Net [27], Test_Meeting [27], SPEECHIO_0, SPEECHIO_1, SPEECHIO_2, SPEECHIO_3, SPEECHIO_4333SPEECHIO_* means SPEECHIO_ASR_ZH0000* [31]. The internal noisy and accent test sets have different acoustic conditions (noisy and accented speech) from the open test sets.
@@ 4.4 Optimal configuration
We also compare our results to vanilla fine-tuned Whisper models and U2++ conformer model [37]. We use the same training set to fine-tune Whisper-large-v2, while the WeNet team[38] fine-tunes Whisper-large-v3 using WenetSpeech (corrected transcript). Additionally, we train a standard U2++ conformer model using the same training data. We initialize the model with parameters from the WeNet team’s publicly available model  [^10] https://github.com/wenet-e2e/wenet/blob/main/docs/pretrained_models.en.md, which is trained solely on WenetSpeech data, and train it for 2 epochs using our training data. The experimental results in Table 6 show that our proposed model performs significantly better than the Whisper series models and U2++ conformer model. This indicates that the performance could not be achieved by the speech foundation model encoder alone, and the LLM helped to improve the performance.
Table 6: Comparison with popular models in terms of CER%($\downarrow$). Whisper-L-v2 and Whisper-L-v3 have both undergone fine-tuning. Here, ’AISL1’ refers to the AISHELL-1 test set, while ’AISL2’ refers to the AISHELL-2 test set. ’Test_N’ and ’Test_M’ represent the Test_Net and Test_Meeting test sets, respectively. ’SIO_’ denotes the SPEECHIO_ test sets.
Dataset Para-former Qwen-Audio Whisper-L-v2 Whisper-L-v3 U2++ Ours
AISL1 1.95 1.30 3.18 5.02 1.44 0.95
AISL2 3.01 3.18 4.49 6.60 3.92 3.50
Tset_N 6.74 9.50 9.08 6.54 8.32 6.06
Test_M 6.97 10.87 9.73 8.96 8.84 6.26
SIO_0 2.55 5.08 3.44 4.05 3.38 3.05
SIO_1 0.49 1.17 1.98 3.17 1.87 1.58
SIO_2 3.23 5.68 5.27 6.25 4.90 3.73
SIO_3 1.13 2.81 4.86 7.09 3.88 2.42
SIO_4 1.33 3.99 3.63 4.29 3.59 3.29